OpenStudy (anonymous):

Simple algebra question (if even that advanced). Expand: (x+ 2) / sqrt(x) I know the answer is sqrt(x) + 2/sqrt(x) but I really have no idea how that answer was achieved. Any help is appreciated.

OpenStudy (hoblos):

(x+ 2) / sqrt(x) = x/sqrtx + 2/sqrtx = sqrtx + 2/sqrtx

OpenStudy (anonymous):

\[\frac{x+2}{\sqrt{x}}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Alright, how the heck is \[x/\sqrt{x} = \sqrt{x}\] ?

OpenStudy (hoblos):

x/sqrtx = x/x^1/2 = x * x^-1/2 = x^1/2 = sqrtx

OpenStudy (anonymous):

Ha, thanks. I mean, I'm taking Calculus II and stuff like this sometimes gets me. Maybe I need to go back to developmental math :)